OpenStudy (anonymous):

A point charge Q1=−2 μC is located at x=0, and a point charge Q2=+8 μC is placed at x=−0.5 m on the x-axis of a cartesian coordinate system.

OpenStudy (anonymous):

 What is E(x) (in N/C) for x=-43.5 m Just want a quick conformation of whether or not it's 29.428 N/C @Mashy

OpenStudy (anonymous):

this is the first homework right?? just check the answer.. duh!!

OpenStudy (anonymous):

i mean in that. you can click on CHECK ANSWER

OpenStudy (anonymous):

I says it's wrong :P I've checked it a million times

OpenStudy (yrelhan4):

whats the question?

OpenStudy (anonymous):

i remember getting the answer right :D

OpenStudy (anonymous):

dang it, mashy :P

OpenStudy (anonymous):

do the calculation man... it needs calci.. i dunno where my calci is..

OpenStudy (anonymous):

wait.. lemme login and see the answers

OpenStudy (anonymous):

haha and i only did the calculation for a couple thousand times so......

OpenStudy (anonymous):

my net is slow..!! it not loading :(

OpenStudy (anonymous):

wait wait.. wait arrr

OpenStudy (anonymous):

the numbers are not the same for everybody. they're randomly generated.

OpenStudy (anonymous):

oh is it.. cause i remember the same numbers for me too :P

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

wait hold on laa!

OpenStudy (anonymous):

thats how TV TALKS :D hahaha

OpenStudy (anonymous):

bahahaaha

OpenStudy (anonymous):

ok the first one right??

OpenStudy (anonymous):

uh huh

OpenStudy (anonymous):

oh yea.. the main question is same.. but the number is differnt lol sorry :D

OpenStudy (anonymous):

lol toldja :P

OpenStudy (anonymous):

good luck :D.. you must be doing something wrong.. show your work here!

OpenStudy (anonymous):

\[\Large9 \times 10^{9} * 2 \times 10^{-6} (\frac{4}{43^{2}}-\frac{1}{43.5^{2}})\] that gives me 29.428

OpenStudy (anonymous):

this is from edx ryt?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yup